OpenStudy (anonymous):

Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2).

OpenStudy (anonymous):

y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1

OpenStudy (anonymous):

First you need to find the slope. And use the given point and the slope to determain the required equation. To find the slope use the formula m1m2=-1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i get it now